OpenStudy (anonymous):

The area of the parallelogram shown is thirty-six millimeters squared. What is the perimeter? Show your work. A parallelogram is shown. The slanted sides are labeled fifteen millimeters. A dotted line extends from the lower right corner of the parallelogram to the right. This line intersects a second dotted line that stretches from the upper right corner of the parallelogram downward, intersecting the first dotted line to form a right angle. This dotted line is labeled 9 millimeters.

OpenStudy (anonymous):

who ever gets this right gets a medal

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

@TylerD

OpenStudy (anonymous):

28

OpenStudy (anonymous):

shoe your work please

OpenStudy (anonymous):

show

OpenStudy (anonymous):

Ok so you know that 10 will be a in the equation so plug it in and solve that way for b. b×10=20mm2 When you solve for b, b will equal 2. so all the sides of the shape equal 13+13+2+2=perimeter. So try solving the last part yourself

OpenStudy (anonymous):

@denisse_stump

OpenStudy (anonymous):

for b which would be the base of the parallelogram. Then I gave you the equation to solve for the perimeter: 13+13+2+2=perimeter

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

but how is it 28

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

ignore 28 sorry

OpenStudy (anonymous):

so its 30

OpenStudy (anonymous):

yes